# Conformal prediction with local weights: randomization enables robust guarantees

### v4

@@ 4 Theoretical guarantees
In this section, we present theoretical coverage guarantees for randomly-localized conformal prediction (RLCP). We will see that these guarantees verify that RLCP achieves the aims outlined earlier in (A1)(A3).
We will begin in Section 4.1 by establishing a key property of the RCLP method in Proposition 1 that underlies in all its coverage guarantees. With this result in place, we then present the marginal coverage guarantee (aim (A1)). Approximate test-conditional coverage (aim (A2)) and robustness to covariate shift (aim (A3)) will be addressed in Sections 4.2 and 4.3, respectively, where we will also explore the relationship between these two problems. RLCP also enjoys a training-conditional coverage guarantee, but since this result is of a different flavor, we defer it to Appendix A.2.
@@ 4.1 Marginal coverage and a key property
Now we return to the RLCP method. The quantile $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ is calculated using a reweighted empirical distribution of the scores. Examining the definition of these weights $\widetilde{w}_{i}$ (defined in Section 3), and recalling the weighted conformal prediction method introduced in Section 1.1, we can see that, conditional on ${\widetilde{X}}_{n+1}$, the RLCP procedure is equivalent to running weighted conformal prediction (WCP) for this covariate shift problem—i.e., when the shifted covariate distribution is given by ${\widetilde{P}}_{X}=P_{X}\circ H(\cdot,{\widetilde{X}}_{n+1})$. The desired coverage result then follows as a direct application of Tibshirani et al., (2019)’s result for weighted conformal prediction, which guarantees coverage at level $1-\alpha$ when weighted conformal prediction is applied to a setting with a (known) covariate shift. ∎
@@ 4.2.1 Example: coverage conditional on a ball
To make the results of Theorem 2 more concrete, let us see what these results imply for a specific setting: assuming ${\mathcal{X}}=\mathbb{R}^{d}$, we will take $B=\mathbb{B}(x_{0},r)=\{x\in\mathbb{R}^{d}:\|x-x_{0}\|_{2}\leq r\}$, the ball of radius $r$ centered at some point $x_{0}$. We will also take $H$ to be the normalized box kernel with bandwidth $h$, $H(x,x^{\prime})=\frac{1}{V_{d}h^{d}}\mathbbm{1}\{\|x-x^{\prime}\|_{2}\leq h\}$, with $2h\leq r$. Applying Theorem 2 to this setting (with $\varepsilon=h$) yields the guarantee
@@
We remark that under some additional smoothness assumptions, we can also derive asymptotic test conditional coverage of the form (2) under this setting—that is, pointwise coverage, which conditions on $X_{n+1}=x_{0}$, rather than $X_{n+1}\approx x_{0}$ as in the result above. We defer the theorem and its proof to the Appendix A.3.
@@ 4.3 Coverage against arbitrary covariate shift
As have been stated before, test-conditional coverage (or approximations of test-conditional coverage, as in Theorem 2) can also be viewed as a special case of coverage against covariate shift settings. To consider a general setting, recall that, for any reweighting function $g:{\mathcal{X}}\rightarrow\mathbb{R}_{\geq 0}$, with ${\mathbb{E}}_{P_{X}}[g(X)]$ positive and finite, the reweighted covariate distribution $P_{X}\circ g$ is defined as in (11). We will consider a setting where training data is drawn from $P=P_{X}\times P_{Y|X}$, while the test point distribution (i.e., the target distribution for which we would like to have predictive coverage) is $(P_{X}\circ g)\times P_{Y|X}$. For example, in a setting where each data point represents a patient and the feature space ${\mathcal{X}}$ includes information about age, $g$ might be a monotone increasing function of age, indicating that older patients comprise a larger fraction of the general population as compared to the training sample; equivalently, older patients are underrepresented in the training sample.
@@
The RLCP method defined in (15) satisfies the following property: for any distribution $P$, and every $g:{\mathcal{X}}\rightarrow\mathbb{R}_{\geq 0}$ with $P_{X}\circ g$ well-defined (i.e., with $0<{\mathbb{E}}_{P_{X}}[g(X)]<\infty$), it holds that
@@ 4.3.1 A closer look at the covariate shift guarantee: proof sketch of Theorem 3
Now we can ask, for what functions $g$ might we expect this total variation distance to be small? If $H$ is a localizing kernel with a fairly small bandwidth, then the reweighted distribution $P_{X}\circ H(\cdot,{\widetilde{X}})$ is likely to place nearly all of its mass in a small neighborhood around ${\widetilde{X}}$. Consequently, if $g$ is Lipschitz (or some similar condition), then $g$ will be approximately constant on the bulk of the distribution $P_{X}\circ H(\cdot,{\widetilde{X}})$, and thus reweighting additionally by $g$ (i.e., computing the new distribution $P_{X}\circ g\circ H(\cdot,{\widetilde{X}})$) will only minimally change this distribution. The localization is very much necessary to this argument: if we ran conformal prediction without localization (i.e., CP rather than one of the LCP methods), following the same arguments would lead us to a term $\textnormal{d}_{\textnormal{TV}}(P_{X},P_{X}\circ g)$, which is likely to be large even if $g$ is Lipschitz (aside from the trivial case where $g$ is a constant function, i.e., $P_{X}\circ g=P_{X}$).
@@
We can see that the conditional variance in the numerator is likely to be small when $H$ is strongly localized (i.e., when the conditional distribution of $X\mid{\widetilde{X}}$ is concentrated in a small neighborhood), while $g$ is Lipschitz, so that $g(X)$ has low variability conditional on ${\widetilde{X}}$. The remaining details, including proving the bound (20) and also establishing the upper bound in the theorem, are deferred to Appendix A.1.
@@ 5.1.1 Univariate setting
In Appendix C.3, in addition to showing plots for all values of the bandwidth $h$ (since Figure 2 only shows results for $h=0.1,0.4$), we also compare to a related method from the recent literature—the work of Gibbs et al., (2023), which we will describe below in Section 6.1.
@@ 5.1.2 Multivariate setting
These weights appear in the calculation of the weighted quantile for both baseLCP and calLCP. To understand how the $w_{i}$’s relate to effective sample size, let’s consider the simple case where $H$ is a box kernel and each $X$ is likely to pick $k$ many neighbors out of a sample of size $n$ with this kernel: then for each $i$, we have (approximately) $k$ many $w_{i}$’s with value (approximately) $1/k$, and the rest zero. In this setting, for each $i$ we have $\sum_{i}w_{i}^{2}\approx 1/k$. Equivalently, we can write
@@
The results, which are displayed in Figure 4, show that calLCP has more disparity in the coverage conditional on $B_{\textnormal{in}}$ versus on $B_{\textnormal{out}}$, as compared to RLCP: aside from the lowest values of dimension $d$, the coverage of RLCP on these subsets are around $1\%$ off from the target level, compared to around $3\%$ for calLCP. For both of these methods, as expected from the theory, the marginal coverage is exactly $1-\alpha$. (For baseLCP, the marginal coverage is quite conservative except at the smallest values of $d$, and hence the conditional coverage levels are high as well.)
@@
The results shown above consider conditional coverage relative to only two specific sets $B$—namely, $B_{\textnormal{in}}$ and $B_{\textnormal{out}}$. For a more complete picture of the performance of the methods, in Appendix C.4, we consider a more general collection of sets $B$, and compare coverage of all these localized conformal methods conditioned on those sets. We will see a similar pattern of performance—divergence of coverage of calLCP from the target level is significantly higher compared to that of RLCP, for higher dimensions $d$.
@@ 5.2 Real data example: predicting age of abalones
Next, we will compare the performance of the local CP methods on a real dataset on abalones (a species of mollusk) (Nash et al.,, 1994). [^5] The data for this experiment was obtained from https://archive.ics.uci.edu/dataset/1/abalone. The objective is to predict the age of abalones based on their physical measurements. Since the economic value of abalone is positively correlated with its age, but determining age precisely requires a costly procedure, accurately predicting age from physical characteristics alone becomes both a valuable and practical problem. This dataset has $4177$ data points, where the response rings determines the age of the abalone, and the features consist of the abalone’s sex (categorized as Male, Female, or Infant), length (the longest shell measurement), diameter (perpendicular to length), height, whole weight, shucked weight (i.e., weight of the meat inside), viscera weight (i.e., gut weight after bleeding) and shell weight (the weight after drying).
@@
$H(x,x^{\prime})=\frac{1}{2h_{\texttt{sex}}\cdot 2h_{\texttt{diameter}}\cdot 2h_{\texttt{height}}\cdot 2h_{\texttt{whole\_weight}}}\cdot\mathbbm{1}\bigg{\{}x_{\texttt{sex}}=x^{\prime}_{\texttt{sex}},|x_{\texttt{length}}-x^{\prime}_{\texttt{length}}|\leq h_{\texttt{length}}\\ \hskip 142.26378pt|x_{\texttt{diameter}}-x^{\prime}_{\texttt{diameter}}|\leq h_{\texttt{diameter}},|x_{\texttt{height}}-x^{\prime}_{\texttt{height}}|\leq h_{\texttt{height}},\\ |x_{\texttt{whole\_weight}}-x^{\prime}_{\texttt{whole\_weight}}|\leq h_{\texttt{whole\_weight}}\bigg{\}}.$
@@
In the right panel of Figure 6, we ask the same question with respect to $X_{\texttt{length}}$. Since this is a continuous covariate, we cannot compute conditional coverage exactly, but instead estimate it with a sliding window whose width is chosen to have $5\%$ mass: that is, we are plotting ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{length}}\in[x-\Delta_{0},x+\Delta_{1})\}$ where, for each $x$, $\Delta_{0}$ and $\Delta_{1}$ are chosen so that $2.5\%$ of the abalone population has length in the range $[x-\Delta_{0},x)$, and same for the range $[x,x+\Delta_{1})$. Here we again see that RLCP shows better conditional coverage compared to calLCP particularly at lower bandwidth $h=0.05$—the estimated conditional coverage for RLCP is approximately constant at the target level of $90\%$ while for calLCP, we see very prominent over- or undercoverage at various values of length.
@@ 6 Discussion and extensions
In this work, we have introduced a novel variant of the localized conformal prediction framework, randomly-localized conformal prediction (RLCP), which attains distribution-free approximate coverage guarantees for relaxed test-conditional coverage and for smooth covariate shift problems while maintaining exact marginal coverage. As explained in Section 4.3.1, all of these coverage results are consequences of a core total variation distance based argument, which could potentially be used to derive other types of coverage guarantees as well. Additionally, RLCP also enjoys a training-conditional coverage guarantee. Our numerical simulations and data experiments advocate the use of RLCP over existing locally-weighted conformal prediction algorithms, due to its better performance at maintaining conditional coverage across a range of settings.
@@ 6.1 Related work
Another type of approach that can be viewed through the lens of Step (b), is the approach of partitioning ${\mathcal{X}}$ into bins, and then computing $\widehat{q}$ separately within each bin; methods of this type have been explored by, e.g., Lei and Wasserman, (2014); Izbicki et al., (2022); LeRoy and Zhao, (2021), and can be related to Step (b) by choosing $H(x,x^{\prime})=\mathbbm{1}\{\textnormal{$x,x^{\prime}$ are in the same bin}\}$. Alternatively, $H$ can be chosen as a nearest-neighbor based kernel (Ghosh et al.,, 2023).
@@
If we take $\cal F$ to be the set of all measurable functions, this would be equivalent to a test-conditional coverage guarantee as in (2)—but as discussed before, is impossible to achieve distribution-free. Gibbs et al., (2023) therefore implement their method with more restricted classes of functions $\cal F$, for instance linear functions or functions from a RKHS, to achieve an approximate notion of local coverage; their method is implemented via solving a quantile regression problem over the class $\mathcal{F}$. As another example, we can choose =$\mathcal{F}=\{\mathbbm{1}\{\cdot\in G\}:G\in\mathcal{G}\}$ for a collection $\mathcal{G}$ of subsets of ${\mathcal{X}}$; this amounts to requiring group-conditional coverage for all groups from the collection $\mathcal{G}$, as has been studied in Jung et al., (2022).
@@ 6.2 Is it possible to derandomize RLCP?
In our discussion in Section 6.1, we showed that a range of methods from the literature can all be viewed through a three-step framework for constructing localized versions of conformal prediction—including our proposed method, RLCP. One way that RLCP differs from the other methods discussed, however, is the fact that randomization is inherent in its construction. In this section, we take a closer look at the question of randomization: is the randomized construction of RLCP necessary for ensuring its favorable theoretical and empirical properties, and does it cause substantial variability in the output in practice?
@@
Figure 7: A visualization of the variability of the RLCP prediction band for a fixed data set, showing the 5th and 95th percentile of $300$ draws of constructing RLCP using the same simulated data set. We also compare to the oracle prediction interval and to the calLCP prediction interval. See Section 6.2 for details.
Hence, depending on the underlying setting, this variability of RLCP can be substantial, and may be a substantial limitation of this method in practice. On the other hand, the randomized construction is at the heart of the ability of RLCP to obtain stronger and cleaner theoretical guarantees relative to existing methods (as we have discussed in Sections 3.1 and 4.3.1), as well as its favorable empirical performance in terms of local coverage.
With this tradeoff in mind, this naturally suggests the following question: is it possible to “derandomize” the RLCP method, perhaps by averaging or aggregating across multiple runs of RLCP on a given data set, to retain its strong theoretical and empirical properties but reduce the issue of variability due to randomization?
6.2.1 Derandomizing RLCP via $m$-randomly-localized CP ($m$-RLCP)

### v5

@@ 4 Theoretical guarantees
In this section, we present theoretical coverage guarantees for RLCP. We will see that these guarantees verify that RLCP achieves the aims outlined earlier in (A1)(A3).
We will begin in Section 4.1 by establishing a key property of the RCLP method in Proposition 1 that underlies in all its coverage guarantees. With this result in place, we then present the marginal coverage guarantee (aim (A1)). Approximate test-conditional coverage (aim (A2)) and robustness to covariate shift (aim (A3)) will be addressed in Sections 4.2 and 4.3, respectively, where we will also explore the relationship between these two problems. RLCP also enjoys a training-conditional coverage guarantee, but since this result is of a different flavour, we defer it to Appendix A.2.
@@ 4.1 Marginal coverage and a key property
Now we return to the RLCP method. The quantile $\widehat{q}_{1-\alpha}(X_{n+1},{\widetilde{X}}_{n+1})$ is calculated using a reweighted empirical distribution of the scores. Examining the definition of these weights $\widetilde{w}_{i}$ (defined in Section 3), and recalling the weighted conformal prediction method introduced in Section 1.1, we can see that, conditional on ${\widetilde{X}}_{n+1}$, the RLCP procedure is equivalent to running weighted conformal prediction (WCP) for this covariate shift problem—i.e., when the shifted covariate distribution is given by ${\widetilde{P}}_{X}=P_{X}\circ H(\cdot,{\widetilde{X}}_{n+1})$. The desired coverage result then follows as a direct application of Tibshirani et al., (2019)’s result for weighted conformal prediction, which guarantees coverage at level $1-\alpha$ when WCP is applied to a setting with a (known) covariate shift. ∎
@@ 4.2.1 Example: coverage conditional on a ball
To make the results of Theorem 2 more concrete, let us see what these results imply for a specific setting: assuming ${\mathcal{X}}=\mathbb{R}^{d}$, we will take $B=\mathbb{B}(x_{0},r)=\{x\in\mathbb{R}^{d}:\|x-x_{0}\|_{2}\leq r\}$, the ball of radius $r$ centred at some point $x_{0}$. We will also take $H$ to be the normalized box kernel with bandwidth $h$, $H(x,x^{\prime})=\frac{1}{V_{d}h^{d}}\mathbbm{1}\{\|x-x^{\prime}\|_{2}\leq h\}$, with $2h\leq r$. Applying Theorem 2 to this setting (with $\varepsilon=h$) yields the guarantee
@@
We remark that under some additional smoothness assumptions, we can also derive asymptotic test-conditional coverage of the form (2) under this setting—that is, pointwise coverage, which conditions on $X_{n+1}=x_{0}$, rather than $X_{n+1}\approx x_{0}$ as in the result above. We defer the theorem and its proof to the Appendix A.3.
@@ 4.3 Coverage against arbitrary covariate shift
As have been stated before, test-conditional coverage (or approximations of test-conditional coverage, as in Theorem 2) can also be viewed as a special case of coverage against covariate shift settings. To consider a general setting, recall that, for any reweighting function $g:{\mathcal{X}}\rightarrow\mathbb{R}_{\geq 0}$, with ${\mathbb{E}}_{P_{X}}[g(X)]$ positive and finite, the reweighted covariate distribution $P_{X}\circ g$ is defined as in (11). We will consider a setting where training data is drawn from $P=P_{X}\times P_{Y|X}$, while the test point distribution (i.e. the target distribution for which we would like to have predictive coverage) is $(P_{X}\circ g)\times P_{Y|X}$. For example, in a setting where each data point represents a patient and the feature space ${\mathcal{X}}$ includes information about age, $g$ might be a monotone increasing function of age, indicating that older patients comprise a larger fraction of the general population as compared to the training sample; equivalently, older patients are underrepresented in the training sample.
@@
The RLCP method defined in (15) satisfies the following property: for any distribution $P$, and every $g:{\mathcal{X}}\rightarrow\mathbb{R}_{\geq 0}$ with $P_{X}\circ g$ well defined (i.e., with $0<{\mathbb{E}}_{P_{X}}[g(X)]<\infty$), it holds that
@@ 4.3.1 A closer look at the covariate shift guarantee: proof sketch of Theorem 3
Now we can ask, for what functions $g$ might we expect this total variation distance to be small? If $H$ is a localizing kernel with a fairly small bandwidth, then the reweighted distribution $P_{X}\circ H(\cdot,{\widetilde{X}})$ is likely to place nearly all of its mass in a small neighbourhood around ${\widetilde{X}}$. Consequently, if $g$ is Lipschitz (or some similar condition), then $g$ will be approximately constant on the bulk of the distribution $P_{X}\circ H(\cdot,{\widetilde{X}})$, and thus reweighting additionally by $g$ (i.e., computing the new distribution $P_{X}\circ g\circ H(\cdot,{\widetilde{X}})$) will only minimally change this distribution. The localization is very much necessary to this argument: if we ran conformal prediction without localization (i.e., CP rather than one of the LCP methods), following the same arguments would lead us to a term $\textnormal{d}_{\textnormal{TV}}(P_{X},P_{X}\circ g)$, which is likely to be large even if $g$ is Lipschitz (aside from the trivial case where $g$ is a constant function, i.e., $P_{X}\circ g=P_{X}$).
@@
We can see that the conditional variance in the numerator is likely to be small when $H$ is strongly localized (i.e., when the conditional distribution of $X\mid{\widetilde{X}}$ is concentrated in a small neighbourhood), while $g$ is Lipschitz, so that $g(X)$ has low variability conditional on ${\widetilde{X}}$. The remaining details, including proving the bound (20) and also establishing the upper bound in the theorem, are deferred to Appendix A.1.
@@ 5.1.1 Univariate setting
In Appendix C.3, in addition to showing plots for all values of the bandwidth $h$ (since Figure 2 only shows results for $h=0.1,0.4$), we also compare with a related method from the recent literature—the work of Gibbs et al., (2023), which we will describe below in Section 6.1.
@@ 5.1.2 Multivariate setting
These weights appear in the calculation of the weighted quantile for both baseLCP and calLCP. To understand how the $w_{i}$’s relate to effective sample size, let us consider the simple case where $H$ is a box kernel and each $X$ is likely to pick $k$ many neighbours out of a sample of size $n$ with this kernel: then for each $i$, we have (approximately) $k$ many $w_{i}$’s with value (approximately) $1/k$, and the rest zero. In this setting, for each $i$ we have $\sum_{i}w_{i}^{2}\approx 1/k$. Equivalently, we can write
@@
The results, which are displayed in Figure 4, show that calLCP has more disparity in the coverage conditional on $B_{\textnormal{in}}$ versus on $B_{\textnormal{out}}$, as compared with RLCP: aside from the lowest values of dimension $d$, the coverage of RLCP on these subsets are around $1\%$ off from the target level, compared with around $3\%$ for calLCP. For both of these methods, as expected from the theory, the marginal coverage is exactly $1-\alpha$. (For baseLCP, the marginal coverage is quite conservative except at the smallest values of $d$, and hence the conditional coverage levels are high as well.)
@@
The results shown above consider conditional coverage relative to only two specific sets $B$—namely, $B_{\textnormal{in}}$ and $B_{\textnormal{out}}$. For a more complete picture of the performance of the methods, in Appendix C.4, we consider a more general collection of sets $B$, and compare coverage of all these localized conformal methods conditioned on those sets. We will see a similar pattern of performance—divergence of coverage of calLCP from the target level is significantly higher compared with that of RLCP, for higher dimensions $d$.
@@ 5.2 Real data example: predicting age of abalones
Next, we will compare the performance of the localized CP methods on a real dataset on abalones (a species of mollusc) (Nash et al.,, 1994). [^5] The data for this experiment were obtained from https://archive.ics.uci.edu/dataset/1/abalone. The objective is to predict the age of abalones based on their physical measurements. Since the economic value of abalone is positively correlated with its age, but determining age precisely requires a costly procedure, accurately predicting age from physical characteristics alone becomes both a valuable and practical problem. This dataset has $4177$ data points, where the response rings determines the age of the abalone, and the features consist of the abalone’s sex (categorized as Male, Female, or Infant), length (the longest shell measurement), diameter (perpendicular to length), height, whole weight, shucked weight (i.e., weight of the meat inside), viscera weight (i.e., gut weight after bleeding) and shell weight (the weight after drying).
@@
$H(x,x^{\prime})$ $=\frac{1}{2h_{\texttt{sex}}\cdot 2h_{\texttt{diameter}}\cdot 2h_{\texttt{height}}\cdot 2h_{\texttt{whole\_weight}}}\cdot$
$\mathbbm{1}\bigg{\{}x_{\texttt{sex}}=x^{\prime}_{\texttt{sex}},|x_{\texttt{length}}-x^{\prime}_{\texttt{length}}|\leq h_{\texttt{length}},|x_{\texttt{diameter}}-x^{\prime}_{\texttt{diameter}}|\leq h_{\texttt{diameter}},$
$|x_{\texttt{height}}-x^{\prime}_{\texttt{height}}|\leq h_{\texttt{height}},|x_{\texttt{whole\_weight}}-x^{\prime}_{\texttt{whole\_weight}}|\leq h_{\texttt{whole\_weight}}\bigg{\}}.$
@@
In the right panel of Figure 6, we ask the same question with respect to $X_{\texttt{length}}$. Since this is a continuous covariate, we cannot compute conditional coverage exactly, but instead estimate it with a sliding window whose width is chosen to have $5\%$ mass: that is, we are plotting ${\mathbb{P}}\{Y_{n+1}\in\widehat{C}_{n}(X_{n+1})\mid(X_{n+1})_{\texttt{length}}\in[x-\Delta_{0},x+\Delta_{1})\}$ where, for each $x$, $\Delta_{0}$ and $\Delta_{1}$ are chosen so that $2.5\%$ of the abalone population has length in the range $[x-\Delta_{0},x)$, and same for the range $[x,x+\Delta_{1})$. Here we again see that RLCP shows better conditional coverage compared with calLCP particularly at lower bandwidth $h=0.05$—the estimated conditional coverage for RLCP is approximately constant at the target level of $90\%$ while for calLCP, we see very prominent over- or undercoverage at various values of length.
@@ 6 Discussion and extensions
In this work, we have introduced a novel variant of the localized conformal prediction framework, RLCP, which attains distribution-free approximate coverage guarantees for relaxed test-conditional coverage and for smooth covariate shift problems while maintaining exact marginal coverage. As explained in Section 4.3.1, all of these coverage results are consequences of a core total variation distance based argument, which could potentially be used to derive other types of coverage guarantees as well. Additionally, RLCP also enjoys a training-conditional coverage guarantee. Our numerical simulations and data experiments advocate the use of RLCP over existing localized CP algorithms, due to its better performance at maintaining conditional coverage across a range of settings.
@@ 6.1 Related work
Another type of approach that can be viewed through the lens of Step (b), is the approach of partitioning ${\mathcal{X}}$ into bins, and then computing $\widehat{q}$ separately within each bin; methods of this type have been explored by, e.g., Lei and Wasserman, (2014); Izbicki et al., (2022); LeRoy and Zhao, (2021), and can be related to Step (b) by choosing $H(x,x^{\prime})=\mathbbm{1}\{\textnormal{$x,x^{\prime}$ are in the same bin}\}$. Alternatively, $H$ can be chosen as a nearest-neighbour based kernel (Ghosh et al.,, 2023).
@@
If we take $\cal F$ to be the set of all measurable functions, this would be equivalent to a test-conditional coverage guarantee as in (2)—but as discussed before, is impossible to achieve distribution-free. Gibbs et al., (2023) therefore implement their method with more restricted classes of functions $\cal F$, for instance linear functions or functions from a Reproducing Kernel Hilbert Space (RKHS), to achieve an approximate notion of local coverage; their method is implemented via solving a quantile regression problem over the class $\mathcal{F}$. As another example, we can choose =$\mathcal{F}=\{\mathbbm{1}\{\cdot\in G\}:G\in\mathcal{G}\}$ for a collection $\mathcal{G}$ of subsets of ${\mathcal{X}}$; this amounts to requiring group-conditional coverage for all groups from the collection $\mathcal{G}$, as has been studied in Jung et al., (2023).
@@ 6.2 Is it possible to derandomize RLCP?
In our discussion in Section 6.1, we showed that a range of methods from the literature can all be viewed through a three-step framework for constructing localized versions of conformal prediction—including our proposed method, RLCP. One way that RLCP differs from the other methods discussed, however, is the fact that randomization is inherent in its construction. In this section, we take a closer look at the question of randomization: is the randomized construction of RLCP necessary for ensuring its favourable theoretical and empirical properties, and does it cause substantial variability in the output in practice?
@@
Figure 7: A visualization of the variability of the RLCP prediction band for a fixed dataset, showing the 5th and 95th percentile of $300$ draws of constructing RLCP using the same simulated dataset. We also compare to the oracle prediction interval and to the calibrated localized conformal prediction interval. See Section 6.2 for details.
Hence, depending on the underlying setting, this variability of RLCP can be substantial, and may be a substantial limitation of this method in practice. On the other hand, the randomized construction is at the heart of the ability of RLCP to obtain stronger and cleaner theoretical guarantees relative to existing methods (as we have discussed in Sections 3.1 and 4.3.1), as well as its favourable empirical performance in terms of local coverage.
With this trade-off in mind, this naturally suggests the following question: is it possible to “derandomize” the RLCP method, perhaps by averaging or aggregating across multiple runs of RLCP on a given dataset, to retain its strong theoretical and empirical properties but reduce the issue of variability due to randomization?
6.2.1 Derandomizing RLCP via $m$-RLCP